# The asymptotic Plancherel formula and Lusztig’s asymptotic algebra for $\tilde{\mathsf{A}}_{n}$

### v1

@@ 6 Maximal length double coset representatives
By [13, Proposition 2.4] it follows that if $w\in\Gamma_{\lambda}$ (respectively $w\in\Gamma_{\lambda}^{-1}$) then $\ell(\mathsf{w}_{\lambda^{\prime}}w)=\ell(w)-\ell(\mathsf{w}_{\lambda^{\prime}})$ (respectively $\ell(w\mathsf{w}_{\lambda^{\prime}})=\ell(w)-\ell(\mathsf{w}_{\lambda^{\prime}})$). In particular, if $w\in\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}$ then $w$ is of maximal length in its double coset $W_{\lambda^{\prime}}wW_{\lambda^{\prime}}$.
@@
For $\gamma\in P^{(\lambda)}$ and $g\in G_{\lambda}$ we have $\mathsf{m}_{g\gamma}=\mathsf{m}_{\gamma}$, and hence
$\{\mathsf{m}_{\gamma}\mid\gamma\in P^{(\lambda)}\}=\{\mathsf{m}_{\gamma}\mid\gamma\in P^{(\lambda)}_{+}\}.$
@@
By [10, Lemma 2.19] we have $\tau_{g\gamma}=g\tau_{\gamma}g^{-1}$ for all $g\in G_{\lambda}$, and hence
$\mathsf{u}_{\lambda}^{-1}\tau_{g\gamma}\mathsf{u}_{\lambda}=(\mathsf{u}_{\lambda}^{-1}g\mathsf{u}_{\lambda})(\mathsf{u}_{\lambda}^{-1}\tau_{\gamma}\mathsf{u}_{\lambda})(\mathsf{u}_{\lambda}^{-1}g^{-1}\mathsf{u}_{\lambda}),$
and by Proposition 2.10 we have $\mathsf{u}_{\lambda}^{-1}g\mathsf{u}_{\lambda}\in W_{\lambda^{\prime}}$. Thus $\mathsf{u}_{\lambda}^{-1}\tau_{g\gamma}\mathsf{u}_{\lambda}\in W_{\lambda^{\prime}}\mathsf{u}_{\lambda}^{-1}\tau_{\gamma}\mathsf{u}_{\lambda}W_{\lambda^{\prime}}$, and hence the result. ∎
@@
Moreover, if $\gamma,\gamma^{\prime}\in P^{(\lambda)}_{+}$ with $\gamma+Q_{\lambda}\preccurlyeq_{\lambda}\gamma^{\prime}+Q_{\lambda}$ then $\ell(\mathsf{m}_{\gamma})\leq\ell(\mathsf{m}_{\gamma^{\prime}})$ with equality if and only if $\gamma=\gamma^{\prime}$.
@@ 7 Lusztig’s asymptotic algebra
In this section we compute the set $\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}$, and describe the matrix algebra $\mathfrak{C}_{\lambda}$, hence giving an explicit description of Lusztig’s asymptotic algebra $\mathcal{J}_{\lambda}$ (recall from Theorem 5.8 that $\mathfrak{C}_{\lambda}\cong\mathcal{J}_{\lambda}$). Section 7.1 provides preliminary observations, in particular showing that each $\lambda$-leading matrix $\mathfrak{c}_{\lambda}(w)$ (with $w\in\Delta_{\lambda}$) has a unique non-zero entry, with the row (respectively column) of this entry indexed by the right (respectively left) cell containing $w$. Section 7.2 gives an explicit description of the commutative ring
$\mathfrak{C}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}=\mathrm{span}_{\mathbb{Z}}\{\mathfrak{c}_{\lambda}(w)\mid w\in\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}\}$
and in Section 7.3 we give our description of $\mathfrak{C}_{\lambda}$.
@@ 7.1 Preliminary observations
Recall that $N_{\lambda}=(n+1)!/(\lambda_{1}!\lambda_{2}!\cdots\lambda_{r(\lambda)}!).$ Then $N_{\lambda}=\dim(\pi_{\lambda})$ (see Section 2.6) and $N_{\lambda}$ also equals the number of right (or left) cells contained in $\Delta_{\lambda}$ by [28, Theorem 14.4.5]. Recall that $\mathsf{B}_{\lambda}=\{\xi_{\lambda}\otimes X_{u}\mid u\in{{}^{\lambda}}W\}$ is a basis of the module $M_{\lambda}$, and thus choosing an order on ${{}^{\lambda}}W$ determines the rows and columns of the matrices $\pi_{\lambda}(h;\mathsf{B}_{\lambda})$.
@@
7.2 The set $\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}$ and the ring $\mathfrak{C}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}$
@@
Note that if $A,B\in\mathfrak{C}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}$ then by Corollary 7.3 we have $A=a(\zeta_{\lambda})E_{\mathsf{u}_{\lambda},\mathsf{u}_{\lambda}}$ and $B=b(\zeta_{\lambda})E_{\mathsf{u}_{\lambda},\mathsf{u}_{\lambda}}$ for some $a(\zeta_{\lambda}),b(\zeta_{\lambda})\in\mathbb{Z}[\zeta_{\lambda}]$ and then the inner product in Theorem 5.6 becomes
@@ 2. $(2)$
The fact that the Schur functions satisfy (1) and (2) is classical, see [23, Proposition 3.4] (and $a_{\gamma,\gamma^{\prime}}=K_{\gamma,\gamma^{\prime}}\geq 0$ are the Kostka numbers). To prove that these elements are unique, suppose that $\mathfrak{s}_{\gamma}^{\prime}(\zeta_{\lambda})$ satisfy (1) and (2), and that $\mathfrak{s}_{\gamma^{\prime}}^{\prime}(\zeta_{\lambda})$ are determined for all $\gamma^{\prime}\prec_{\lambda}\gamma$. By (1) and (2) it follows that $\{\mathfrak{s}^{\prime}_{\gamma^{\prime}}(\zeta_{\lambda})\mid\gamma^{\prime}\prec_{\lambda}\gamma\}$ is an orthonormal basis for the subspace of $G_{\lambda}$-invariant functions spanned by $\{\mathsf{m}_{\gamma^{\prime}}(\zeta_{\lambda})\mid\gamma^{\prime}\prec_{\lambda}\gamma\}$. Thus $\mathfrak{s}_{\gamma}(\zeta_{\lambda})=\mathfrak{e}_{\gamma}(\zeta_{\lambda})+\sum_{\gamma^{\prime}\prec_{\lambda}\gamma}b_{\gamma,\gamma^{\prime}}\mathfrak{s}_{\gamma^{\prime}}(\zeta_{\lambda})$ for some integers $b_{\gamma,\gamma^{\prime}}$, and these integers are uniquely determined by $b_{\gamma,\gamma^{\prime}}=-\langle\mathfrak{e}_{\gamma}(\zeta_{\lambda}),\mathfrak{s}^{\prime}_{\gamma^{\prime}}(\zeta_{\lambda})\rangle_{\lambda}^{\infty}$, using (2). Hence the result. ∎
@@
    To check that the next $\ell(\mathsf{u}_{\lambda}^{-1}\tau_{\gamma}\mathsf{u}_{\lambda})$ steps can be taken to be crossings we must show that this part of the path remains in $\mathcal{A}_{\lambda}$ (and hence there are no forced bounces). Note that the starting alcove of this part of the path is $\mathsf{u}_{\lambda}\mathcal{A}_{0}$ (which lies in $\mathcal{A}_{\lambda}$) and the end alcove is $\mathsf{u}_{\lambda}(\mathsf{u}_{\lambda}^{-1}\tau_{\gamma}\mathsf{u}_{\lambda})\mathcal{A}_{0}=\tau_{\gamma}\mathsf{u}_{\lambda}\mathcal{A}_{0}$. Thus by [10, Theorem 2.9] the end alcove lies in $\mathcal{A}_{\lambda}$, and since the path is of reduced type, and $\mathcal{A}_{\lambda}$ is convex (being an intersection of half-spaces) the entire path from $\mathsf{u}_{\lambda}\mathcal{A}_{0}$ to $\tau_{\lambda}\mathsf{u}_{\lambda}\mathcal{A}_{0}$ lies in $\mathcal{A}_{\lambda}$ (see [1, Proposition 3.94]).
@@ 3. $(3)$
    To show that the final $\ell(y)$ steps are positive folds, note first that if $s\in J_{\lambda^{\prime}}$ then $\tau_{\lambda}\mathsf{u}_{\lambda}s\mathcal{A}_{0}\subseteq\mathcal{A}_{\lambda}$ (since $\mathsf{u}_{\lambda}s\in{{}^{\lambda}}W$, and apply [10, Theorem 2.9]). Moreover, since $\tau_{\gamma}\mathsf{u}_{\lambda}=t_{\gamma}\mathsf{y}_{\gamma}\mathsf{u}_{\lambda}$ and $\ell(\mathsf{y}_{\gamma}\mathsf{u}_{\lambda}s)=\ell(\mathsf{y}_{\gamma}\mathsf{u}_{\lambda})+1$ we have that $\tau_{\gamma}\mathsf{u}_{\lambda}\mathcal{A}_{0}$ lies on the positive side of the hyperplane separating $\tau_{\gamma}\mathsf{u}_{\lambda}\mathcal{A}_{0}$ from $\tau_{\gamma}\mathsf{u}_{\lambda}s\mathcal{A}_{0}$, and hence the result.
@@
Consider a path $p\in\mathcal{P}_{\lambda}(\vec{\mathsf{m}}_{\gamma},\mathsf{u}_{\lambda})_{\mathsf{u}_{\lambda}}$ with $\mathrm{wt}(p)=\gamma^{\prime}$. Since $\theta^{\lambda}(p)=\mathsf{u}_{\lambda}$ [10, Corollary 2.11] gives $\mathrm{end}(p)=\tau_{\gamma^{\prime}}\mathsf{u}_{\lambda}$. Let $N=f(p)+b(p)$ be the total number of folds and bounces in $p$. If $\vec{\mathsf{m}}_{\gamma}=s_{1}s_{2}\cdots s_{\ell}\pi$ then, since $p$ starts at $\mathsf{u}_{\lambda}$, we have $\mathrm{end}(p)=\mathsf{u}_{\lambda}s_{1}\cdots\hat{s}_{i_{1}}\cdots\hat{s}_{i_{N}}\cdots s_{\ell}\pi,$ where the folds and bounces of $p$ occur at the indices $i_{k}$ with $1\leq i_{1}<i_{2}<\cdots<i_{N}\leq\ell$ (and $\hat{s}_{i_{k}}$ denotes omission of the generator in the expression). It follows that $\ell(\mathsf{u}_{\lambda}^{-1}\tau_{\gamma^{\prime}}\mathsf{u}_{\lambda})\leq\ell(\mathsf{m}_{\gamma})-N,$ and thus $N\leq\ell(\mathsf{m}_{\gamma})-\ell(\mathsf{u}_{\lambda}^{-1}\tau_{\gamma^{\prime}}\mathsf{u}_{\lambda})$. Since $\ell(\mathsf{m}_{\gamma^{\prime}})=\ell(\mathsf{u}_{\lambda}^{-1}\tau_{\gamma^{\prime}}\mathsf{u}_{\lambda})+\ell(\mathsf{w}_{\lambda^{\prime}})$ we have
@@
The following theorem is one of the main theorem of this paper. It gives a $\lambda$-relative version of [15, Proposition 8.6] (see also [23, Theorem 2.22(b)]). Recall that
$\mathfrak{C}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}=\mathrm{span}_{\mathbb{Z}}\{\mathfrak{c}_{\lambda}(w)\mid w\in\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}\}.$
By Theorem 5.8 we have
$\mathfrak{C}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}\cong\mathcal{J}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}.$
@@
$\mathfrak{C}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}\to\mathbb{Z}[\zeta_{\lambda}]^{G_{\lambda}}\quad\text{with}\quad\mathfrak{c}_{\lambda}(\mathsf{m}_{\gamma})\mapsto\mathfrak{s}_{\gamma}(\zeta_{\lambda})$
@@
We can now complete the proof of Lemma 3.11 (and hence the $\lambda$-relative Satake Isomorphism).

### v2

@@ 6 Maximal length double coset representatives
By [16, Proposition 2.4] it follows that if $w\in\Gamma_{\lambda}$ (respectively $w\in\Gamma_{\lambda}^{-1}$) then $\ell(\mathsf{w}_{\lambda^{\prime}}w)=\ell(w)-\ell(\mathsf{w}_{\lambda^{\prime}})$ (respectively $\ell(w\mathsf{w}_{\lambda^{\prime}})=\ell(w)-\ell(\mathsf{w}_{\lambda^{\prime}})$). In particular, if $w\in\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}$ then $w$ is of maximal length in its double coset $W_{\lambda^{\prime}}wW_{\lambda^{\prime}}$.
@@
If $\gamma\in P^{(\lambda)}$ and $g\in G_{\lambda}$ then $\mathsf{m}_{g\gamma}=\mathsf{m}_{\gamma}$.
@@
By [13, Lemma 2.19] we have $\tau_{g\gamma}=g\tau_{\gamma}g^{-1}$ for all $g\in G_{\lambda}$, and hence
$\mathsf{u}_{\lambda}^{-1}\tau_{g\gamma}\mathsf{u}_{\lambda}=(\mathsf{u}_{\lambda}^{-1}g\mathsf{u}_{\lambda})(\mathsf{u}_{\lambda}^{-1}\tau_{\gamma}\mathsf{u}_{\lambda})(\mathsf{u}_{\lambda}^{-1}g^{-1}\mathsf{u}_{\lambda}).$
By Proposition 2.10 we have $\mathsf{u}_{\lambda}^{-1}g\mathsf{u}_{\lambda}\in W_{\lambda^{\prime}}$, so $\mathsf{u}_{\lambda}^{-1}\tau_{g\gamma}\mathsf{u}_{\lambda}\in W_{\lambda^{\prime}}\mathsf{u}_{\lambda}^{-1}\tau_{\gamma}\mathsf{u}_{\lambda}W_{\lambda^{\prime}}$, hence the result. ∎
@@ 7 Lusztig’s asymptotic algebra
In this section we compute the set $\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}$, and describe the matrix algebra $\mathfrak{C}_{\lambda}$, hence giving an explicit description of Lusztig’s asymptotic algebra $\mathcal{J}_{\lambda}$ (recall from Theorem 5.11 that $\mathfrak{C}_{\lambda}\cong\mathcal{J}_{\lambda}$). Section 7.1 provides preliminary observations, in particular showing that each $\lambda$-leading matrix $\mathfrak{c}_{\lambda}(w)$ (with $w\in\Delta_{\lambda}$) has a unique non-zero entry, with the row (respectively column) of this entry indexed by the right (respectively left) cell containing $w$.
Let
$\mathcal{J}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}=\mathrm{span}_{\mathbb{Z}}\{\mathsf{t}_{w}\mid w\in\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}\}\quad\text{and}\quad\mathfrak{C}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}=\mathrm{span}_{\mathbb{Z}}\{\mathfrak{c}_{\lambda}(w)\mid w\in\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}\}.$
Since $\Gamma_{\lambda}^{-1}$ is a left cell $\mathcal{J}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}$ is a ring, and by Theorem 5.11 we have $\mathcal{J}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}\cong\mathfrak{C}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}$. In Section 7.2 we study the ring $\mathfrak{C}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}$, and as a consequence show that $\mathcal{J}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}$ is a commutative ring isomorphic to the ring $\mathbb{Z}[\zeta_{\lambda}]^{G_{\lambda}}$ of $G_{\lambda}$-symmetric functions. This recovers the main result of Xi [34, Theorems 8.2.1 and 8.4.5]. Finally, in Section 7.3 we give our description of $\mathfrak{C}_{\lambda}$, and hence of $\mathcal{J}_{\lambda}$ and the asymptotic algebra $\mathcal{J}=\bigoplus_{\lambda\vdash n+1}\mathcal{J}_{\lambda}$.
@@ 7.1 Preliminary observations
Recall that $N_{\lambda}=(n+1)!/(\lambda_{1}!\lambda_{2}!\cdots\lambda_{r(\lambda)}!).$ Then $N_{\lambda}=\dim(\pi_{\lambda})$ (see Section 2.6) and $N_{\lambda}$ also equals the number of right (or left) cells contained in $\Delta_{\lambda}$ by [31, Theorem 14.4.5]. Recall that $\mathsf{B}_{\lambda}=\{\xi_{\lambda}\otimes X_{u}\mid u\in{{}^{\lambda}}W\}$ is a basis of the module $M_{\lambda}$, and thus choosing an order on ${{}^{\lambda}}W$ determines the rows and columns of the matrices $\pi_{\lambda}(h;\mathsf{B}_{\lambda})$.
@@
7.2 The set $\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}$ and the ring $\mathcal{J}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}\cong\mathfrak{C}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}$
@@
Note that if $A,B\in\mathcal{J}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}$ then by Corollary 7.3 we have $\pi_{\lambda}^{\infty}(A)=a(\zeta_{\lambda})E_{\mathsf{u}_{\lambda},\mathsf{u}_{\lambda}}$ and $\pi_{\lambda}^{\infty}(B)=b(\zeta_{\lambda})E_{\mathsf{u}_{\lambda},\mathsf{u}_{\lambda}}$ for some $a(\zeta_{\lambda}),b(\zeta_{\lambda})\in\mathbb{Z}[\zeta_{\lambda}]$ and then the inner product in Theorem 5.7 becomes
@@ 2. $(2)$
The fact that the Schur functions satisfy (1) and (2) is classical, see [26, Proposition 3.4] (and $a_{\gamma,\gamma^{\prime}}=K_{\gamma,\gamma^{\prime}}\geq 0$ are the Kostka numbers). To prove that these elements are unique, suppose that $\mathfrak{s}_{\gamma}^{\prime}(\zeta_{\lambda})$ satisfy (1) and (2), and that $\mathfrak{s}_{\gamma^{\prime}}^{\prime}(\zeta_{\lambda})$ are determined for all $\gamma^{\prime}\prec_{\lambda}\gamma$. By (1) and (2) it follows that $\{\mathfrak{s}^{\prime}_{\gamma^{\prime}}(\zeta_{\lambda})\mid\gamma^{\prime}\prec_{\lambda}\gamma\}$ is an orthonormal basis for the subspace of $G_{\lambda}$-invariant functions spanned by $\{\mathsf{m}_{\gamma^{\prime}}(\zeta_{\lambda})\mid\gamma^{\prime}\prec_{\lambda}\gamma\}$. Thus $\mathfrak{s}_{\gamma}(\zeta_{\lambda})=\mathfrak{e}_{\gamma}(\zeta_{\lambda})+\sum_{\gamma^{\prime}\prec_{\lambda}\gamma}b_{\gamma,\gamma^{\prime}}\mathfrak{s}_{\gamma^{\prime}}(\zeta_{\lambda})$ for some integers $b_{\gamma,\gamma^{\prime}}$, and these integers are uniquely determined by $b_{\gamma,\gamma^{\prime}}=-\langle\mathfrak{e}_{\gamma}(\zeta_{\lambda}),\mathfrak{s}^{\prime}_{\gamma^{\prime}}(\zeta_{\lambda})\rangle_{\lambda}^{\infty}$, using (2). Hence the result. ∎
@@
    To check that the next $\ell(\mathsf{u}_{\lambda}^{-1}\tau_{\gamma}\mathsf{u}_{\lambda})$ steps can be taken to be crossings we must show that this part of the path remains in $\mathcal{A}_{\lambda}$ (and hence there are no forced bounces). Note that the starting alcove of this part of the path is $\mathsf{u}_{\lambda}\mathcal{A}_{0}$ (which lies in $\mathcal{A}_{\lambda}$) and the end alcove is $\mathsf{u}_{\lambda}(\mathsf{u}_{\lambda}^{-1}\tau_{\gamma}\mathsf{u}_{\lambda})\mathcal{A}_{0}=\tau_{\gamma}\mathsf{u}_{\lambda}\mathcal{A}_{0}$. Thus by [13, Theorem 2.9] the end alcove lies in $\mathcal{A}_{\lambda}$, and since the path is of reduced type, and $\mathcal{A}_{\lambda}$ is convex (being an intersection of half-spaces) the entire path from $\mathsf{u}_{\lambda}\mathcal{A}_{0}$ to $\tau_{\lambda}\mathsf{u}_{\lambda}\mathcal{A}_{0}$ lies in $\mathcal{A}_{\lambda}$ (see [1, Proposition 3.94]).
@@ 3. $(3)$
    To show that the final $\ell(y)$ steps are positive folds, note first that if $s\in J_{\lambda^{\prime}}$ then $\tau_{\lambda}\mathsf{u}_{\lambda}s\mathcal{A}_{0}\subseteq\mathcal{A}_{\lambda}$ (since $\mathsf{u}_{\lambda}s\in{{}^{\lambda}}W$, and apply [13, Theorem 2.9]). Moreover, since $\tau_{\gamma}\mathsf{u}_{\lambda}=t_{\gamma}\mathsf{y}_{\gamma}\mathsf{u}_{\lambda}$ and $\ell(\mathsf{y}_{\gamma}\mathsf{u}_{\lambda}s)=\ell(\mathsf{y}_{\gamma}\mathsf{u}_{\lambda})+1$ we have that $\tau_{\gamma}\mathsf{u}_{\lambda}\mathcal{A}_{0}$ lies on the positive side of the hyperplane separating $\tau_{\gamma}\mathsf{u}_{\lambda}\mathcal{A}_{0}$ from $\tau_{\gamma}\mathsf{u}_{\lambda}s\mathcal{A}_{0}$, and hence the result.
@@
Consider a path $p\in\mathcal{P}_{\lambda}(\vec{\mathsf{m}}_{\gamma},\mathsf{u}_{\lambda})_{\mathsf{u}_{\lambda}}$ with $\mathrm{wt}(p)=\gamma^{\prime}$. Since $\theta^{\lambda}(p)=\mathsf{u}_{\lambda}$ [13, Corollary 2.11] gives $\mathrm{end}(p)=\tau_{\gamma^{\prime}}\mathsf{u}_{\lambda}$. Let $N=f(p)+b(p)$ be the total number of folds and bounces in $p$. If $\vec{\mathsf{m}}_{\gamma}=s_{1}s_{2}\cdots s_{\ell}\pi$ then, since $p$ starts at $\mathsf{u}_{\lambda}$, we have $\mathrm{end}(p)=\mathsf{u}_{\lambda}s_{1}\cdots\hat{s}_{i_{1}}\cdots\hat{s}_{i_{N}}\cdots s_{\ell}\pi,$ where the folds and bounces of $p$ occur at the indices $i_{k}$ with $1\leq i_{1}<i_{2}<\cdots<i_{N}\leq\ell$ (and $\hat{s}_{i_{k}}$ denotes omission of the generator in the expression). It follows that $\ell(\mathsf{u}_{\lambda}^{-1}\tau_{\gamma^{\prime}}\mathsf{u}_{\lambda})\leq\ell(\mathsf{m}_{\gamma})-N,$ and thus $N\leq\ell(\mathsf{m}_{\gamma})-\ell(\mathsf{u}_{\lambda}^{-1}\tau_{\gamma^{\prime}}\mathsf{u}_{\lambda})$. Since $\ell(\mathsf{m}_{\gamma^{\prime}})=\ell(\mathsf{u}_{\lambda}^{-1}\tau_{\gamma^{\prime}}\mathsf{u}_{\lambda})+\ell(\mathsf{w}_{\lambda^{\prime}})$ we have
@@
The following theorem is one of the main theorems of this paper. It gives a $\lambda$-relative version of [18, Proposition 8.6] (see also [26, Theorem 2.22(b)]), and recovers the main result of Xi [34]. However we note that our description of $\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}$ and $\mathcal{J}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}$ is rather different to the descriptions given by Xi [34] and Kim and Pylyavskyy [17]. In particular, our description of $\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}$ is in terms of the geometry of the fundamental $\lambda$-alcove $\mathcal{A}_{\lambda}$, and our description of the ring $\mathcal{J}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}$ is in terms of symmetric functions and alcove paths, rather than representation rings of algebraic groups.
@@
$\mathcal{J}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}\to\mathbb{Z}[\zeta_{\lambda}]^{G_{\lambda}}\quad\text{with}\quad\mathsf{t}_{\mathsf{m}_{\gamma}}\mapsto\mathfrak{s}_{\gamma}(\zeta_{\lambda})$
@@
We can now complete the proof of Lemma 3.11.
@@
Remark 7.7.
The description of $\mathcal{J}_{\Gamma_{\lambda}\cap\Gamma_{\lambda}^{-1}}$ in [34, Theorem 8.4.5] is in terms of representation rings. To make the translation with our description given in Theorem 7.6, for each $\lambda\vdash n+1$ let $u_{\lambda}$ be a unipotent element of $\mathsf{SL}_{n+1}(\mathbb{C})$ with Jordan blocks given by the partition $\lambda$. Let $F_{\lambda}$ be the maximal reductive subgroup of the centraliser in $\mathsf{SL}_{n+1}(\mathbb{C})$ of $u_{\lambda}$. Then $\mathbb{Z}[\zeta_{\lambda}]^{G_{\lambda}}$ is isomorphic to the representation ring of $F_{\lambda}$.
We also note that Xi works with the “affine symmetric group” $W_{\mathrm{aff}}\rtimes\mathbb{Z}$, and our extended affine Weyl group $\widetilde{W}$ is a quotient of this group. Thus Xi’s results apply to $\mathsf{GL}_{n+1}(\mathbb{C})$. The setup of our paper can readily be adapted to work with the affine symmetric group and its associated extended affine Hecke algebra, however we have chosen to present our work directly in terms of the extended affine Weyl group $P\rtimes W$ (and hence $\mathsf{SL}_{n+1}(\mathbb{C})$) as this setting more naturally extends to other types (where an “extended affine Weyl group” is typically taken to be a group of the form $L\rtimes W$ with $L$ a lattice with $Q\leq L\leq P$).